OpenStudy (anonymous):

Find 5x - 3y - z if x = -2, y = 2, and z = -3

OpenStudy (anonymous):

Easy! Just substitute everything in. It would look like this: 5(-2)-3(2)-(-3) 5 x -2 = -10 -3 x 2 = -6 -3 so then... -10-6-3 now you just combine them all. so your answer would be -19.